OpenStudy (anonymous):

Solve 11.3^(2x – 1) = 15.7

OpenStudy (anonymous):

(2x-1)log11.3 = log15.7 2x-1 = log15.7 / log11.3 2x = (log15.7 / log11.3) + 1 x = [(log15.7 / log11.3) + 1] / 2